OpenStudy (anonymous):

About 11% of the population is math phobic. If two people are randomly selected, what is the probability that both are math phobic? What is the probability that at least one is math phobic?

OpenStudy (anonymous):

Two independent events. Probability of both = P(A)*P(B)

OpenStudy (anonymous):

P(both phobic) = .11 * .11 = 0.121 or 12.1 %

OpenStudy (anonymous):

P(at least one) = 1-P(neither)

OpenStudy (amistre64):

\[\frac{\binom{11}{2}}{\binom{100}{2}}\] maybe?

OpenStudy (zarkon):

amistre64...that would assume you are only looking at a population of 100 people.

OpenStudy (amistre64):

right, i was wondering if i could generalize the 11% like that

OpenStudy (amistre64):

whats the earths population?

OpenStudy (anonymous):

No, it's better to just think of them as independent events with a set likelihood each time.

OpenStudy (anonymous):

Ahahaha.

OpenStudy (amistre64):

the bigger my numbers the more it stabalizes at .012 ... ;)

OpenStudy (anonymous):

Makin' the problem so complicated lols. But it's more accurate too, I suppose.

OpenStudy (zarkon):

we assume an infinite population...using 1,000,000 or infinity will yield almost the same result

OpenStudy (amistre64):

.11*.11 = .0121 rught?

OpenStudy (zarkon):

yes